# From Wavefunctional Entanglement to Entangled Wavefunctional Degrees of Freedom

### v2

@@ Abstract
The question of whether entanglement between photons is equivalent to entanglement between their characteristic field modes—specifically, the single-particle wavefunctions that are composed and superposed to describe identical particles—is a key, open problem concerning multi-partite optical degrees of freedom, and has profound implications for topics ranging from quantum foundations to quantum computation. Here, I offer a fresh, deeper, physical insight into this subtle, albeit enduring, issue by describing a situation in which entangling interactions between optical modes—namely, the wavefunctions—can be distilled into genuine entanglement between the physical properties of the photons—which are the wavefunctional degrees of freedom. This theoretical observation also highlights the salience of the measurement context—especially, of clearly disambiguating between the choice of the quantum subsystem and the decision to measure an observable along a particular axis of measurement—while quantifying and transforming quantum optical entanglement. This understanding might be applied to formulate a new class of protocols for distilling quantum magic from contextually and nonlocally entangled photons within inseparable field modes.
@@ I Introduction
A more careful reading of their paper reveals that E.P.R. devised entanglement as a mechanism to predict the value of a physical quantity of a quantum system, say, system I with “certainty” [1]—that is, with probability equal to unity, like in a classical, realist theory—without in any way “disturbing” [1]—that is, altering the physical state of—the system. The actual, direct measurement were to be carried out on the system’s entangled, spatially separated counterpart, say, system II. Crucially, E.P.R. implicitly assumed that the principle of classical locality should ensure that such an indirect measurement on system II would not “disturb” system I [2, 43].
Guided by Bohm’s reformulation of the E.P.R. paradox [44] in terms of discrete-valued, dichotomous observables, namely, spin angular momenta, Bell correctly realized that the above-described conflict is actually one between the irrevocable indeterminism of quantum mechanics and E.P.R.’s presupposed, cherished notion of local, classical realism [45, *bell2004speakablepaper2]. Simply put, situations, such as the E.P.R.B. (E.P.R.-Bohm) entanglement in which there are perfect correlations—or anti-correlations—between the outcomes of measurements performed upon two causally disconnected, space-like separated particles suggest that there are two opposing descriptions of quantum reality, namely, either: physical reality is nonlocal, measurement outcomes are inescapably probabilistic, and the wavefunction description of physical reality is complete; or physical reality is local, measurement outcomes are predetermined by hidden elements of reality, and therefore the wavefunction description of physical reality is incomplete, which is the crux of the E.P.R. argument [47, *werner2014comment, *wiseman2014two].
@@
II Motivation for this work
The quantum mechanical systems of choice in a majority of the above-described experiments were optical, which raised an accompanying, profound question: (Qu1) Is entanglement between modes of light equivalent to entanglement between individual particles of light? (see, for example, Refs. [82, 83, 84] for a detailed discussion on the history of, and related work on this key problem.) Resolution of this quantum foundational issue also has significant implications for quantum information processing [85] and quantum-enhanced sensing [86] with highly entangled states of light.
@@ II Motivation for this work
Consequently, one might pose the question, which seems both fundamentally, as well as practically important: (Qu3) Is it possible to transfer the inter-modal entanglement—which is relatively simpler to achieve—to inter-particle entanglement—which is particularly challenging to realize—as deterministically as possible? The purpose of this paper is to show that, indeed, wavefunctional 111Throughout this paper, the term wavefunctional is used as the adjectival form of wavefunction, and is unrelated to a wave functional that plays the role of a wavefunction in a quantized theory of wave fields [120]. entanglement—that is, entanglement between mathematically inseparable quantum wavefunctions or optical field modes—can be transformed into direct entangling interactions between optical photons that are described by such modes. An additional step that entails making a measurement on an ancillary photon appears to be irreplaceable. Schematically, Fig. 1 summarizes the problem that this paper addresses, whereas Fig. 2 depicts a proposed solution.
@@
It is important to note that pairs of indistinguishable optical photons—whose angles of linear polarization are entangled—are routinely produced by spontaneous parametric down-conversion (SPDC)-based setups [90]. However, although these sources of entangled photon pairs can be, for example, cascaded [91, 92] to generate three-photon entangled states [92, 93], such highly optically non-linear procedures require complex experimental arrangements [94, 95], and are plagued by low conversion efficiencies, thereby imposing severe requirement on the pump laser. In this paper, I propose an alternative strategy that entails creating mode-mode entangled states—where the modes contain up to photons—and converting such mode-mode entangled states to genuine photon-photon entangled states.
III Disambiguating between choices of subsystems and of observables
@@ IV The Equivalence between Wavefunctional and Inter-particle Entanglement
With the above caveats in mind, we are now ready to address (Qu3). The principal findings are summarized in the following principle:
@@
The Equivalence between Wavefunctional and Inter-particle Entanglement Principle. For a quantum optical system having spatial modes, and operating with input, unentangled photons, the wavefunctional entanglement—that is, the entanglement between the single-particle quantum wavefunctions, or modes—can be transformed into entangling interactions between the photons that are described by such modes.
@@
While I have given a somewhat more formal proof in Appendix C, here, I will outline a simple argument that provides a deeper physical intuition into attacking this formidable problem. Figure 2 schematically summarizes this physical argument. Inspired by Ref. [87], I will model these spatial, bosonic modes as bound states, as opposed to, for example, propagating Gaussian wavepackets. I am specifically assuming that such modes can be mapped onto the bound energy eigenmodes of a simple harmonic oscillator, and of its anharmonically perturbed analogues; a quantum mechanical particle can always, meaningfully, said to be placed in a bound state.
@@
Consider the following input state that describes two modes, which are in a photon occupation number entangled relationship with each other:
@@
where and represent the two optical modes that are mapped onto the two oscillator energy eigenmodes, and and label the two photons that are supposed to be distinguishable. The symbol, indicates that the subsystems on the two sides of this symbol are different; for example, the entangled subsystems on the left (right) hand side above are modes (particles). Assuming identical photons would have implied: (up to normalization constants)—a phenomenon known as entanglement due to symmetrization [101, 82]. As this case is trivial and not useful for applications, I am considering distinguishable photons occupying distinguishable modes, with additional photonic degrees of freedom—for instance, angular momenta—ensuring that the overall state is symmetrized. To be as consistent as possible, in this simple Gedankenexperiment, with how optical fields are quantized in the second quantization approach, let us assume that the single-photon states map only onto the single-quantum oscillator eigenmodes. The zero- and two-photon states are assumed to map onto distinct oscillator modes, which, presumably, can be adiabatically eliminated from this problem, because of their distinct eigenenergies. Consequently, as depicted in Fig. 2(b), let us focus on the middle term, which corresponds to two photons occupying two distinct modes.
Let us now deploy a quantum mechanical random process that we will utilize to create macroscopic quantum superpositions. To accomplish this, we can introduce an “ancillary” quantum mechanical system, namely, a photon (photon 3 in Fig. 2) propagating through the Young’s double-slit experimental setup [see Fig. 2(a)]. As is well known, the single photon—as it transits through this apparatus—interferes with itself, and, consequently, is described by the superposition: of having gone through the top and bottom slits. Moreover, assume the screen to be photo-conducting, and without a lower half—so as to further emphasize that, unlike in the classical case, the position on the screen, for example, the upper half, where the photon is detected is uncorrelated with the slit of entry, for example, the top slit. The screen performs a measurement of the position of the photon. Now—assuming ideal detection and quantum efficiencies—the overall state of the detector and amplifier system, which has access only to photons impinging on the upper half, can be written as the superposition: of having registered and not registered the single photon as a single photo-electron. Assuming that the regions of photo-detection and collection of photo-electrons are sufficiently macroscopic, the probability amplitude, akin to the expression for the transition probability amplitude of a macroscopically delocalized detector atom for this diffraction experiment [102]. describes the spectral representation of the electromagnetic field propagating from the source to the detector.
@@
The probabilistically ejected photo-electron is amplified, thereby producing a photo-current that drives an anharmonic potential controller, which controls the anharmonicity parameter, ; for example, no photo-current corresponds to the initially harmonic potential. Consequently, the mode wavefunctions are now described as macroscopic superpositions of the energy eigenfunctions of a harmonic oscillator, and their perturbed counterparts corresponding to an anharmonically altered oscillator.
@@
Any realization of this Gedankenexperiment will have two requisites, namely: (A) an “ancillary” quantum mechanical subsystem has to be coupled to the entangling system; and (B) a quantum measurement has to be made on this subsystem—as are required for entangling photons that have never interacted in the past [103]. Hence, I have argued that mode-entangled states can be transformed into particle-entangled states by anharmonically and controllably—by a quantum mechanical random process—deforming an initially harmonic potential.

### v3

@@ Abstract
The question of whether entanglement between photons is equivalent to entanglement between their characteristic field modes—specifically, the single-particle wavefunctions that are composed and superposed to describe particles in such modes—is a key, open problem concerning multi-partite optical degrees of freedom, and has profound implications for topics ranging from quantum foundations to quantum computation. Here, I offer a fresh, deeper, physical insight into this subtle, albeit enduring, issue by describing a situation in which entangling interactions between optical modes—namely, the wavefunctions—can be distilled into genuine entanglement between the physical, observable properties of the photons—which are the wavefunctional degrees of freedom. This theoretical observation also highlights the salience of the measurement context—especially, of clearly disambiguating between the choice of the quantum subsystem and the decision to measure an observable along a particular axis of measurement—while quantifying and transforming quantum optical entanglement. This theoretical observation might be applied to formulate a new class of protocols for performing quantum information tasks, using entangled photons within inseparable field modes.
@@ I Introduction
A more careful reading of their paper reveals that E.P.R. devised entanglement as a mechanism to predict the value of a physical quantity of a quantum system, say, system I with certainty [1]—that is, with probability equal to unity, like in a classical, realist theory—without in any way disturbing [1]—that is, altering the physical state of—the system. The actual, direct measurement were to be carried out on the system’s entangled, spatially separated counterpart, say, system II. Crucially, E.P.R. implicitly assumed that the principle of classical locality should ensure that such an indirect measurement on system II would not disturb system I [2, 43].
Guided by Bohm’s reformulation of the E.P.R. paradox [44] in terms of discrete-valued, dichotomous observables, namely, spin angular momenta, Bell correctly realized that the above-described conflict is actually one between the irrevocable indeterminism of quantum mechanics and E.P.R.’s presupposed, cherished notion of local, classical realism [45, *bell2004speakablepaper2]. Simply put, situations, such as the E.P.R.B. (E.P.R.-Bohm) entanglement in which there are perfect correlations—or anti-correlations—between the outcomes of measurements performed upon two causally disconnected, space-like separated particles suggest that there are two opposing descriptions of quantum reality, namely, either: physical reality is nonlocal, measurement outcomes are inescapably probabilistic, and the wavefunctional description of physical reality is complete; or physical reality is local, measurement outcomes are predetermined by hidden elements of reality, and, therefore, the wavefunctional description of physical reality is incomplete, which is the crux of the E.P.R. argument [47, *werner2014comment, *wiseman2014two].
@@
The quantum mechanical systems of choice in a majority of the above-described experiments were optical, which raised an accompanying, profound question: (Qu1) Is entanglement between modes of light equivalent to entanglement between individual particles of light? Resolution of this quantum foundational issue also has significant implications for quantum information processing [82] and quantum-enhanced sensing [83] with highly entangled states of light. While Refs. [84, 85, 86] comprehensively survey the relevant theoretical and experimental advancements in understanding the similarities and the differences between mode-mode entanglement and particle-particle entanglement, the remainder of this introduction offers a selection of the most illuminating and interesting aspects—of such advances—so as to motivate the importance of elucidating whether entanglement between optical photons is equivalent to entanglement between their characteristic optical field modes.
The following discussion is inspired, in particular, by the systematic survey offered in the End Matter of Ref. [86]. As is well known and is well understood, it is essential to clearly and unambiguously identify the relevant quantum subsystems—as well as the salient quantum observables that are entangled across such quantum subsystems—so as to properly and rigorously characterize and quantify quantum mechanical entanglement. For the case of composite, multi-partite, quantum optical systems, one can consider either the modes of light, or the particles of light, as the individual quantum subsystems. In the literature, usually, distinguishable modes are regarded as the more suitable candidate, since such modes have all the essential characteristics of well-defined quantum subsystems [84, 87, 88, 89, 90, 91, 92, 65, 93]. For instance, modes can be distinguishable; modes are directly, physically accessible and are amenable to measurements [94]; and, finally, modes can exist as separate, spatially distinct systems that can be prepared in quantum states that correspond to that individual system alone. In contrast, considering mathematically labeled, individual, indistinguishable particles as quantum subsystems appears to be more of a mathematical construct, as compared with a physically reasonable approach, because such particles do not fulfill the above-described criteria, for well-defined quantum subsystems. Specifically, Peres has emphasized that the properly exchange-symmetrized wavefunctions of two identical particles are always entangled, if one views the individual, indistinguishable particles as the subsystems. This notion is often referred to as entanglement due to symmetrization [95]. This entanglement due to symmetrization, however, is usually neither useful nor utilizable for applications in quantum information science and engineering, with the possible exception of the scheme presented in Ref. [96]; see, also, the discussion about this issue in Sec. IV and Fig. 2.
In general, however, algorithms, protocols, and schemes, in quantum information science and engineering, are designed and implemented, such that they operate with and act upon quantum subsystems, for instance, qubits that are encoded from quantum mechanical particles, such as neutral atoms, trapped ions, atomic nuclei, polar molecules, as well as individual photons and electrons. The present work is, therefore, motivated by the reasonable hope and goal of identifying and constructing genuinely entangled states of optical modes, as well as genuinely entangled states of photons—that circumvents the issue of entanglement due to symmetrization—and ascertaining if and how the former can be transformed into the latter.
@@
Consequently, one might pose the question, which seems both fundamentally, as well as practically important: (Qu3) Is it possible to transfer the inter-modal entanglement—which is relatively simpler to achieve—to inter-particle entanglement—which is particularly challenging to realize—as deterministically as possible? The purpose of this paper is to show that, indeed, wavefunctional 111Throughout this paper, the term wavefunctional is used as the adjectival form of wavefunction, and is unrelated to a wave functional that plays the role of a wavefunction in a quantized theory of wave fields [139]. entanglement—that is, entanglement between mathematically inseparable quantum wavefunctions or optical field modes—can be transformed into direct entangling interactions between optical photons that are described by such modes. An additional step that entails making a measurement on an ancillary photon appears to be irreplaceable. Schematically, Fig. 1 summarizes the problem that this paper addresses.
@@
It is important to note that pairs of indistinguishable optical photons—whose angles of linear polarization are entangled—are routinely produced by spontaneous parametric down-conversion (SPDC)-based setups [100]. However, although these sources of entangled photon pairs can be, for example, cascaded [101, 102] to generate three-photon entangled states [102, 103], such highly optically non-linear procedures require complex experimental arrangements [104, 105], and are plagued by low conversion efficiencies, thereby imposing severe requirement on the pump laser. In this paper, I propose an alternative strategy that entails creating mode-mode entangled states—where the modes contain up to photons—and converting such mode-mode entangled states to genuine photon-photon entangled states. Therefore, when viewed from an alternative perspective, the aim of this work can also be construed as suggesting experimentally feasible ways of transforming the entanglement between two arbitrary, bosonic, optical modes into entanglement between their constituent, characteristic, spatial modes.
I will now briefly describe how the remainder of this paper is arranged. Section II highlights the importance of disambiguating between the choices of the quantum subsystems and the pertinent quantum observables, while considering quantum states of entangled photons, and of optical modes that are in photon occupation number entangled relationships with each other. In Sec. III, I describe and distinguish between the two fundamental forms of quantum optical entanglement, namely, wavefunctional entanglement and entangled wavefunctional degrees of freedom. While entanglement between optical modes is a particular case of the former, entanglement between particles of light is a specialized case of the latter. This section also explicitly defines the term, particles in this context.
In Sec. IV, an enduring, foundational question in quantum optics, specifically, whether entanglement between optical photons—namely, the particles of light—is equivalent to entanglement between their characteristic field modes, is considered and investigated. In this section, I approach this problem, in a novel manner, by describing a situation, specifically, a Gedankenexperiment in which the entangling interactions between optical modes are distilled into genuine entanglement between the wavefunctional—that is, physical—degrees of freedom of the photons. This distillation is accomplished by making measurements on an ancillary photon. This measurement process—specifically, the detection of the ancillary photon—controls an anharmonic potential that is essential for creating the final entangled state of photons. This Gedankenexperiment can also be construed as an entangling scheme or protocol. Whereas, in Sec. V, I compare and contrast the above scheme with standard entanglement swapping procedures, in Sec. VI, I describe several consequences of non-ideal photo-detection efficiencies, and how such limitations might be overcome. In Sec. VII, the above Gedankenexperiment is simplified and is analyzed, using the notion of quantum pre-measurement. Finally, in Sec. VIII, I summarize the main conclusions of this paper, and suggest future avenues of fundamental research based on this work.
II Disambiguating between choices of subsystems and of observables
@@
III Distinguishing between Wavefunctional Entanglement and Entangled Wavefunctional Degrees of Freedom
The purpose of this section is to introduce and clearly define two key terms, namely, wavefunctional entanglement, and entangled wavefunctional degrees of freedom. One of the central purposes of this paper is to effectively distinguish between these two fundamental forms of quantum entanglement, especially in the context of quantum optical systems and observables. It is important that these two concepts are adequately and sharply defined, as they will play essential, defining roles in the main results of this paper.
The term, wavefunctional entanglement conveys a form of quantum entanglement that is directly evident in, and arises from, the mathematical representation of many-particle wavefunctions as non-factorable, or inseparable, combinations of their constituent, single-particle wavefunctions. In particular, conventional, mode-mode entanglement—for instance, entanglement between bosonic, optical, field modes—is a specific kind of wavefunctional entanglement, since, typically, these modes are modeled using single-particle wavefunctions that are suitably composed and superposed, so as to describe particles within such modes. Wavefunctional entanglement, consequently, is a more widely generalizable notion that attempts to indicate a fundamental form of quantum entanglement that originates from the mathematical inseparability of constituent, single-particle wavefunctions.
The term, entangled wavefunctional degrees of freedom, in contrast, signifies genuine, non-local—and, possibly, in certain kinds of situations involving multiple commuting observables, contextual—entanglement between the physical properties of two or more particles. These wavefunctional degrees of freedom are nothing but the physical degrees of freedom that are intrinsic to the particle, such as the photon, and that are fully specified by the wavefunctions, or, more specifically, the state vectors of the particles, together with the appropriate Hermitian operators for the observables. Said differently, these entangled wavefunctional degrees of freedom are characteristic of genuinely entangled quantum states of particles, or of particle-particle entanglement. Importantly, the meaning and interpretation of the term, particles, which plays a central role in characterizing such entangled wavefunctional degrees of freedom, will depend on the measurement context and the specific details of the entangling protocol. See, for example, the end of Sec. IV for a discussion of the meaning of particles in the context of the entangling scheme that is introduced in that section. Simply put, whereas mode-mode entanglement signifies entanglement between wave-like quantum subsystems, with continuous-valued quantum observables, particle-particle entanglement describes entanglement between particle-like quantum subsystems, admitting of discrete-valued quantum observables.
In particular, with regard to mode-mode entanglement, in this paper, I will specifically consider entanglement between spatial optical modes. To develop a deeper physical intuition into the relationship between mode-mode entanglement and particle-particle entanglement, one might imagine the former as being analogous to entanglement between two distinct, separated regions of space, whereas the latter can be thought of as entanglement between particles in such spaces. Spatial delocalization of the particles of light between the two quantum subsystems—namely, the two spatial optical modes, or the two regions of space—appears to be essential for the equivalence between mode-mode entanglement and particle-particle entanglement.
@@ IV The Equivalence between Wavefunctional and Inter-particle Entanglement
With the above caveats and definitions in mind, we are now ready to address (Qu3). The principal findings, which constitute the central, original results of this paper, are summarized in the following principle:
@@
The Equivalence between Wavefunctional and Inter-particle Entanglement Principle. For a quantum optical system having spatial, optical modes, and operating with input, unentangled photons, the wavefunctional entanglement—that is, the entanglement between the single-particle quantum wavefunctions, or modes—can be transformed into entangling interactions between the photons that are described by such modes.
@@
While I have given a somewhat more formal proof in Appendix C, as well as a somewhat more rigorous analysis in Sec. VII, here, I will outline a simple argument that provides a deeper physical intuition into attacking this formidable problem. Figure 3 schematically summarizes this physical argument. Inspired by Ref. [97], I will model these spatial, bosonic modes as bound states, as opposed to, for example, propagating Gaussian wavepackets. I am specifically assuming that such modes can be mapped onto the bound energy eigenmodes of a simple harmonic oscillator, and of its anharmonically perturbed analogues; a quantum mechanical particle can always be, meaningfully, said to be placed in a bound state.
@@
Consider, for example, the following input state that describes two modes, which are in a photon occupation number entangled relationship with each other:
where and represent the two optical modes that are mapped onto the two oscillator energy eigenmodes, and and label the two photons that are supposed to be distinguishable, by virtue of an extrinsic degree of freedom. The symbol, indicates that the subsystems on the two sides of this symbol are different; for example, the entangled subsystems on the left (right) hand side above are modes (particles).
Notice that if photons 1 and 2 were assumed to be identical, then the middle term of Eq. (3) would have transformed as: (up to normalization constants). This appearance of entanglement, as a consequence of the photons’ being identical bosons and conforming to exchange symmetry, is a phenomenon known as entanglement due to symmetrization [95, 84]. I have described this phenomenon in Fig. 2(a), after mapping the spatial, optical modes onto the bound energy eigenmodes of a quantum potential, specifically, a simple harmonic potential. Consequently, one might represent the above state as: , where and represent two distinct energy eigenfunctions of the potential, and and indicate the spatial coordinates of photons and , respectively. It is well known that this appearance of entanglement is trivial, and not useful for applications, such as quantum information processing with distinguishable qubits. For instance, two-qubit, entangling gate protocols operating with degrees of freedom—which are used to encode the qubits—require that such degrees of freedom are macroscopically distinguishable, so as to distinguish between the control and the target qubits.
In contrast, consider the situation, as shown in Fig. 2(b), of two distinguishable photons, and , which are delocalized across a harmonic potential and an anharmonic potential. If we assume that two distinct, complete and orthonormal, sets of spatial, optical modes: and are mapped onto the energy eigenfunctions of the simple harmonic potential and an anharmonic potential, which are, and , respectively, then such an arrangement can support genuinely entangled states of two photons, for instance, ; see, for example, below for a scheme that describes how such a state can be prepared. Crucially—in this example, as well as in the discussion below—I am considering distinguishable photons that occupy distinguishable modes. Such photons are assumed to be distinguishable, due to an additional, external, photonic degree of freedom, for instance, orbital angular momentum, or an extrinsic, translational degree of freedom.
@@
To be as consistent as possible, in this simple Gedankenexperiment, with how optical fields are quantized in the second quantization approach, let us assume that the single-photon states, such as and , map only onto the single-quantum oscillator eigenmodes. The zero- and two-photon states, such as , , , and , are assumed to map onto distinct oscillator modes, which, presumably, can be adiabatically eliminated from this problem, because of their distinct eigenenergies. Consequently, as depicted in Fig. 3(b), let us focus on the middle term, which corresponds to two photons occupying two distinct modes.
The formulation of the following scheme is partly inspired by the analysis of the Young’s double-slit diffraction experiment, with single quanta, in the classic textbook of Schiff [111]. First, however, I wish to provide a key clarification. Typically, detectors and amplifiers are macroscopic, classical objects, and, consequently, quantum mechanical wavefunctions cannot be assigned to such objects, especially, in accordance with the standard, statistical interpretation of quantum mechanics, specifically, the Copenhagen interpretation of quantum mechanics according to Bohr. Nonetheless, in the scheme below, inspired by Schiff’s analysis, the detector is taken to be an atom, initially in its ground state, that can be treated as a two-level system. Consequently, the detector, in this particular situation, is a microscopic, measuring apparatus. Also, let us suppose the amplifier to be a linear, phase-invariant—also known as phase-insensitive or phase-preserving—amplifier, whose operation can be fully described by the principles of quantum mechanics. It is reasonable to suppose that a single, macroscopic quantum wavefunction can be assigned to such a quantum optical, phase-coherent amplifier, akin to how a single, macroscopic quantum wavefunction can be assigned to a Bose-Einstein condensate of atoms that is configured to function as a matter wave amplifier, or a matter wave interferometer. Therefore, wavefunctions will be assigned to amplifiers and detectors, in the scheme below. Second, a more rigorous analysis of this scheme, without assigning wavefunctions to amplifiers and detectors, and through the lens of quantum pre-measurement followed by detection, is given in Sec. VII.
Let us now deploy a quantum mechanical random process that we will utilize to create macroscopic quantum superpositions. To accomplish this creation, we can introduce an ancillary quantum mechanical system, specifically, photon 3, in Fig. 3, propagating through the Young’s double-slit experimental setup [see Fig. 3(a)]. As is well known, the single photon—as it transits through this apparatus—interferes with itself, and, consequently, is described by the superposition: of having gone through the top and the bottom slits.
Assume, furthermore, that the screen is photo-conducting, and is without a lower half—so as to further emphasize that, unlike in the classical case, the position on the screen, for example, the upper half, where the photon is detected, is uncorrelated with the slit of entry, for example, the top slit. Three clarifying comments are in order here, so as to give a more thorough justification for this design choice. First, exactly identical results will be obtained, if only the lower half, as opposed to the upper half, were present, as a consequence of the symmetry of the setup; see, for example, the dotted, horizontal line in Fig. 3(a). Second, the lack of correlation between the detection location and the slit originates, fundamentally, from the interference of the single photon with itself—as it transits through the double-slit apparatus—and, therefore, should be independent of the spatial extent of the screen, for instance, if only the upper half were present, or if both halves were utilized. In essence, this lack of correlation stems from the fact that photons behave as quantum mechanical probability amplitudes, as opposed to classical particles. Third, and most interestingly, the above design choice was inspired by a subtle, albeit striking, difference between the predictions of Bohmian mechanics and quantum mechanics, for the outcome of the double-slit experiment. The de Broglie-Bohm pilot wave theory, which is the paradigmatic form of self-consistent, deterministic, hidden variable theories, postulates the existence of hypothetical point corpuscles 222Harvey R. Brown and David Wallace, in [140], makes an excellent argument as to why the particles in the quantum theory of de Broglie and Bohm should be called as hypothetical point corpuscles. over and above the usual wave function of quantum mechanics. Notice, again, the axis of symmetry in Fig. 3(a). No photon, according to such a theory, going through the top slit crosses this axis of symmetry, and the same is true for the photon going through the bottom slit. Consequently, a photon that lands on the top half of the screen must have come through the top slit; likewise, a photon that lands on the bottom half of the screen must have come through the bottom slit [113, 114]. To summarize, the above design choice utilizes not only the usual lack of correlation—as compared with classical mechanics, in which the detection location, for particles, is highly correlated with the slit—but also an enhanced lack of correlation over and above Bohmian mechanics, in which the half-plane of detection is correlated with the slit of entry. I would like to point out that the removal of the lower half of the screen, so as to emphasize, or, possibly, enhance this lack of correlation, is neither an essential nor a crucial feature of this scheme, but is merely an added feature.
The screen performs a measurement of the position of the photon. Now—assuming ideal detection and quantum efficiencies—the overall state of the detector and amplifier system, which has access only to photons impinging on the upper half, can be written as the superposition: of having registered and not registered the single photon as a single photo-electron. Notice that the detection of photon 3, or lack thereof, is described as having created a macroscopic superposition of the detector-amplifier system. Single-photon detection, however, typically collapses the wavefunction. Nevertheless, the formation and the maintenance of the macroscopic quantum superposition can be explained, using the notion of quantum pre-measurement. Actually, the above quantum state is intended to convey a genuine, coherent, linear superposition of the event in which photon 3 is detected, and the one in which photon 3 is not detected; specifically, a state that describes the process of quantum pre-measurement. As is well known, the interaction of a photon with a classical, macroscopic detector involves a fundamentally irreversible and decoherent process, which gives rise to, for example, a classical mixture of an event corresponding to detection and an event corresponding to an absence of detection. Conversely, however, the interaction of a photon with a single, detector atom, which can be treated as a two-level quantum system, gives rise to a coherent linear superposition of an excited atom (and no photon) and an unexcited atom (and a single photon). This state describes the situation preceding the actual measurement upon, or the actual detection of, the photon. Therefore, the words, measurement and detection, in this particular section, have to be understood in the light of the above clarification.
Assuming that the regions of photo-detection and collection of photo-electrons are sufficiently macroscopic, the probability amplitude, akin to the expression for the transition probability amplitude of a macroscopically delocalized detector atom for this diffraction experiment [111]. describes the spectral representation of the electromagnetic field propagating from the source to the detector.
@@
The probabilistically ejected photo-electron is amplified by the phase-coherent amplifier, A, thereby producing a quantum photo-current that drives an anharmonic potential controller, which controls the anharmonicity parameter, ; for example, no photo-current corresponds to the initially harmonic potential. Consequently, the modal wavefunctions are now described as macroscopic superpositions of the energy eigenfunctions of a harmonic oscillator, and their perturbed counterparts corresponding to an anharmonically altered oscillator.
@@
Physically, the above state transformation makes use of the well-known notion of amplification of quantum mechanical superpositions. For instance, if a macroscopic superposition of the amplifier-detector system were not formed—such as in the case of a classical mixture of an event corresponding to detection of photon 3, and an event corresponding to an absence of detection of photon 3—then the subsequently collapsed term, would have corresponded to the initial term, , whereas the collapsed term, would have generated .
To model the effects of the anharmonic potential controller, one could write down the following two-initial-state, two-particle, effective Hamiltonian:
where the initially harmonic energy eigenstates, and are supposed to smoothly transform into the anharmonically perturbed energy eigenstates, and , and vice versa; and the particles, 1 and 2 are assumed to make no quantum jumps, either between and , or between and . To elucidate how such a Hamiltonian might reliably generate the desired entangled state, consider the following simple, illustrative example. For small values of the perturbation parameter, and short times, , the initial state, will evolve as:
@@
Any realization of this Gedankenexperiment will have two requisites, namely: (A) an ancillary quantum mechanical subsystem has to be coupled to the entangling system; and (B) a quantum measurement has to be made on this subsystem—as are required for entangling photons that have never interacted in the past [115]. Hence, I have argued that mode-mode entangled states can be transformed into particle-particle entangled states by anharmonically and controllably—by a quantum mechanical random process—deforming an initially harmonic potential. In particular, in the context of the present scheme, the term, “particles” signifies the occupancies of the spatial optical modes. Consequently, one could also interpret this scheme as a quantum measurement-based entangling protocol that transforms the entanglement between two arbitrary, bosonic, optical modes into entanglement between their constituent, characteristic, spatial modes.
Notably, this final photon entanglement differs from the path entanglement that is generated, using traditional methods through mode interference, since the paths of these photons—corresponding to, for example, distinct interferometric paths within a multi-mode interferometer—can be described and modeled as genuine spatial optical modes. I would also like to emphasize that the physical degrees of freedom—that encapsulate the entanglement following the conversion process—are the photon occupation numbers within a conserved total photon occupation number subspace. I speculate, however, that it is certainly possible to propose and implement similar schemes and protocols that convert the wavefunctional entanglement into other physical degrees of freedom that are intrinsic to the photon, and that are fully specified by the wavefunctions, or, more specifically, the state vectors of the photons, together with the appropriate Hermitian operator for the observable. Proposing such schemes and protocols, however, is beyond the scope of this paper.
The current scheme is based on two strong assumptions, both of whose experimental feasibilities, in the context of current technological capabilities, impose severe limitations on the practical utility of this scheme. First, as has been discussed earlier, the photons are assumed to be distinguishable due to an extrinsic degree of freedom, for instance, orbital angular momentum, or some other rotational state. In particular, experiments have already shown the possibility of accessing and manipulating the orbital angular momenta of optical photons [116, 117]. Second, the efficiency of detecting the ancillary photon 3 is assumed to be unity, and the formation of the macroscopic superposition of the detector-amplifier system is assumed to be ideal, without any accompanying errors. Section VI discusses how this issue can be addressed for non-ideal photodetection efficiencies in the context of realistic, present-day detectors.